OpenStudy (anonymous):

The are of a rectangle piece of land is 280 square meters. If the length of the land was 5 meters less and the width was 5 meters more, the shape of the land would be a square. Write an equation to find the width (x) of the land. (Please show steps so I understand, I always medal the best answer and fan everybody.)

OpenStudy (anonymous):

This should help you: http://openstudy.com/study#/updates/52d08031e4b01e5fc1deafb9

OpenStudy (epoweritheta):

let x,y be the length nad breadth respectively . then by given note x-5=y+5 ie x=y+10 but xy=280 ; (y+10)y=280 ; solve for y , say what u get

OpenStudy (anonymous):

I got it, but thank you both for helping me :)